# video-SALMONN 2: Captioning-Enhanced Audio-Visual Large Language Models

### v1

@@ 3.2 Training Strategies
Audio-visual SFT is applied after aligning the audio modality, utilising annotated video data to train the model for synchronised and integrated audio-visual understanding based on interleaved audio and visual token sequences. During this SFT stage, the model is optimised using cross-entropy loss, with video descriptions serving as ground-truth labels that encapsulate both audio and visual information. To further enhance the backbone LLM’s capability in processing paired audio-visual token sequences, a LoRA [36] adapter is used and jointly trained. Additionally, the audio aligner is trained to map the audio encoder’s output to the input representation space of the LLM, ensuring seamless interpretation of audio tokens by the backbone model.
@@
Figure 2: An overview of the training process including audio modality alignment, audio-visual SFT and MrDPO. LoRA is introduced during the audio-visual SFT stage. In each round of the MrDPO training, a new LoRA proxy is added while the old LoRA is merged into the LLM, so the model always contains only one activated LoRA.
3.3 Multi-round DPO
@@ 3.3.1 A Metric based on Atomic Events
Before DPO training, ground-truth video captions are first decomposed into atomic events using a powerful text LLM, such as GPT-4o, to provide references for evaluating the captions generated by the model. Next, we apply the nucleus sampling method [52] to generate caption pairs from the model’s output distribution for the input video.
@@ 3.3.2 Multi-round DPO with LoRA Proxy
    where $\mathbf{W}_{t}$ and $\mathbf{W}_{t-1}$ are the weight parameters to adapt in $\Lambda_{t}$ and $\Lambda_{t-1}$, $\alpha$ is the scaling factor of LoRA, $r$ is the rank of LoRA, $d$ is the dimension of $\mathbf{W}_{t-1}$. $\mathbf{A}_{t-1}\in\mathcal{R}^{d\times r}$ and $\mathbf{B}_{t-1}\in\mathcal{R}^{r\times d}$ are the low-rank matrix parameters of LoRA in the previous round $t-1$, and $\mathbf{W}\in\mathcal{R}^{d\times d}$ is the parameter of LLM backbone.
@@
    where $\mathbf{\pi}_{\theta}=\{\Lambda_{t},\tilde{\Delta}_{t}\}$ and $\pi_{\text{ref}}=\Lambda_{t}$ represent the policy and reference models for round $t$ respectively, $\sigma$ is the sigmoid function, and $\beta$ is a hyper-parameter controlling the deviation from $\pi_{\text{ref}}$. Variables in the first term follow the definitions in the standard DPO loss [35]. In the second term corresponding to cross-entropy learning towards the ground-truth video captions, $\lambda$ is the weight of the second term in the overall loss, $\mathcal{D}_{\text{gt}}$ denotes the SFT training dataset, and $(\mathbf{x},\mathbf{y}_{\text{gt}})$ corresponds to a video and its paired ground-truth text description. $\mathbf{y}_{\text{win}}$ and $\mathbf{y}_{\text{lose}}$ are preferred and dispreferred video captions generated by $\Lambda_{t}$ for $\mathbf{x}$, which are judged using a text LLM based on the atomic-event-based metric proposed in Section 3.3.1. Each mini-batch of training samples is randomly selected from $\mathcal{D}_{\text{gt}}$.
@@ 4.1 Model Specifications
video-SALMONN 2 is built on an internally trained high-performance visual LLM, which is further fine-tuned on LLaVA-OneVision-7B [15]. The model processes video frames at a per-second frame rate of 1 (i.e., $\phi=1$), and can handle up to 110 frames. For videos longer than 110 seconds, 110 frames are uniformly sampled from the video.
For the audio branch, we use the Whisper-Large-v3 encoder [53] as the audio encoder, and a window-level Q-Former [54] with a window length of 0.5 seconds as the audio aligner, producing a total of 60 audio tokens for a 30-second input. The Whisper encoder has a maximum processing duration of $t_{\text{max}}=30$ seconds. The rank $r$ and scaling factor $\alpha$ of LoRA are set to 128 and 2.0, respectively. During training, the visual encoder, the audio encoder, and LLM remain frozen.
@@ 4.2 Data and Training Specifications
In the audio modality alignment stage, LibriSpeech-960h [55] and AudioCaps [56] are used to train the audio aligner. LibriSpeech-960h is utilized for speech recognition training, while AudioCaps is employed for audio captioning training. In the audio-visual SFT stage, experiments are performed using about 13k videos with rich audio information from LLaVA-Video-178k [16] and FineVideo [57]. Both video captioning and video QA are trained during audio-visual SFT.
@@
In the MrDPO stage, before each training round, the model generates a pair of captions for each video in the SFT dataset. To determine whether a caption pair is suitable for DPO, we evaluate the missing information rate and hallucination rate using GPT-3.5. It is also feasible to use a smaller model for evaluation, as shown in Appendix A. If deemed suitable, one caption is selected as the preferred sample, while the other is rejected. The selection criteria for each round are detailed in Appendix B. We use a learning rate of $2\times 10^{-5}$ in the first two DPO rounds, $1\times 10^{-5}$ for the third round, and $2\times 10^{-6}$ afterwards. We explore the values of $\lambda$ in Eqn. (3) in Appendix C and finally set it to 0.1 throughout the entire MrDPO stage. Since only video captioning is trained during MrDPO, the video QA performance of the model will slightly decline. Therefore, SFT data for video QA will be added to the training process in the final round of MrDPO to enhance the video QA performance of the final model.
@@
For the test dataset, we curated a video captioning benchmark to evaluate the event missing rate (Miss) and hallucination rate (Hall). Details of the test data can be found in Appendix D. The benchmark consists of 483 carefully selected videos, each labelled with complete audio-visual captions by human annotators. Atomic events for the test dataset were initially obtained using GPT-4o and then manually refined. The benchmark is open-sourced [^1] https://huggingface.co/datasets/videoSALMONN2/video-SALMONN_2_testset. GPT-3.5 is used to evaluate the event missing and hallucination rates of the generated captions. Details are shown in Appendix E.
@@
Regarding training settings, the resource consumption for each training step is shown in Table 1. In our experiments, we run a total of six gDPO rounds in MrDPO.
@@
5 Experimental Results
5.1 Overall Results
The results of different models are presented in Table 2. In terms of video captioning, video-SALMONN 2 outperforms other models in both information missing and hallucination rates on our caption benchmark. Among existing open-source multimodal LLMs, few can provide detailed and accurate video descriptions, whether purely visual models like Qwen2.5-VL [17], InternVideo 2.5 [58] and VideoLLaMA 3 [19], or audio-visual models like video-SALMONN [30], VideoLLaMA 2 [29] and Qwen2.5-Omni [45]. The caption quality of Qwen2.5-VL is the best among open-source models. Notably, some open-source models, such as Video-LLaMA 3, tend to generate shorter captions, leading to relatively high information missing rates but low hallucination rates. Existing open-source audio-visual LLMs like video-SALMONN and VideoLLaMA 2 fail to generate high-quality video captions. video-SALMONN has a very high total error rate, while VideoLLaMA 2’s captions are too short to contain all the details in the video. Qwen2.5-Omni can perceive both audio and visual information simultaneously, but still mainly focuses on visual information, often omitting human speech in audio and other sound events in captions. Among commercial models, GPT-4o and Gemini 1.5 Pro are capable of generating higher-quality and more detailed captions compared to current open-source models. However, the visual-only version of GPT-4o lacks audio comprehension, leading to a higher rate of missed events. Meanwhile, Gemini’s understanding of visual content is somewhat limited, causing both models to exhibit varying degrees of information omission and hallucination.
Trained based on a visual base model with average performance, video-SALMONN 2 can ultimately achieve the SOTA performance in video captioning, even surpassing commercial models like GPT-4o and Gemini-1.5-Pro. On our caption benchmark, video-SALMONN 2 has reduced the event missing rate by more than 13%, the event hallucination rate by more than 14%, and the total error rate by nearly 28%, compared with the visual base model. The models’ performance on our caption benchmark is highly correlated with their performance on the "detailed" subset of the VDC caption benchmark [41], since both benchmarks evaluate the completeness of the model’s captions. In the VDC benchmark, which evaluates video captions from different perspectives, video-SALMONN 2 also achieves the SOTA results in the “detailed" subset among similar-size models, indicating the effectiveness of our training methods. We also perform Elo ranking with human evaluators inspired by [41], which validates the alignment of our proposed metrics with human perception and highlights the strong video captioning capabilities of video-SALMONN 2. Details are provided in Appendix F.
Besides, as an audio-visual LLM, video-SALMONN 2 retains strong visual understanding capabilities and performs well on various video benchmarks. In benchmarks like Video-MME [59], where audio information is taken into account in benchmark design, video-SALMONN 2 significantly outperforms the visual base model and is also superior to other open-source models. On purely visual benchmarks like NeXT-QA [60] and MLVU [61], video-SALMONN 2 also obtains competitive results.
Table 2: Results of our benchmark for video captioning evaluation. “A” and “V” refer to the audio and visual modalities, respectively. The event missing rate (Miss), hallucination rate (Hall), and total error rate (Total = Miss + Hall) are assessed for the captions. Results of some QA benchmarks like Video-MME (VMME), NeXT-QA (NQA) and MLVU are listed here as well.
Model Modality Our Caption Benchmark VDC Detailed VMME NQA MLVU

### v3

@@ 3.2 Multi-Stage Training for Audio–Visual Captioning
Audio-visual SFT follows audio modality alignment and uses annotated videos to train the model for synchronised, integrated audio–visual understanding over interleaved audio and visual token sequences. We optimise with cross-entropy loss, using video descriptions as ground-truth labels that capture both modalities. To strengthen the backbone LLM’s handling of paired sequences, we attach a LoRA adapter and train it jointly (while keeping the encoders frozen). In parallel, the audio aligner is further trained to project audio-encoder outputs into the LLM’s input space, ensuring seamless interpretation of audio tokens by the backbone.
@@
Figure 2: An overview of the training process, including audio modality alignment, audio-visual SFT, and MrDPO. LoRA is introduced during the audio-visual SFT stage. In each round of the MrDPO training, a new LoRA proxy is added while the old LoRA is merged into the LLM, so the model always contains only one activated LoRA. The MrDPO-trained model can be applied to annotate video captions.
3.3 Reinforcement Learning with Multi-round DPO
@@ 3.3.1 A Metric based on Atomic Events
Before DPO training, ground-truth video captions are first decomposed into atomic events using a powerful text LLM, such as GPT-4o, to provide references for evaluating the captions generated by the model. Next, we apply the nucleus sampling method (Holtzman et al., 2020) to generate caption pairs from the model’s output distribution for the input video.
@@ 3.3.2 Multi-round DPO with LoRA Proxy
    where $\mathbf{W}_{t}$ and $\mathbf{W}_{t-1}$ are the weight parameters to adapt in $\Lambda_{t}$ and $\Lambda_{t-1}$, $\alpha$ is the scaling factor of LoRA, $r$ is the rank of LoRA, $d$ is the dimension of $\mathbf{W}_{t-1}$. $\mathbf{A}_{t-1}\in\mathcal{R}^{d\times r}$ and $\mathbf{B}_{t-1}\in\mathcal{R}^{r\times d}$ are the low-rank matrix parameters of LoRA in the previous round t1t-1, and $\mathbf{W}\in\mathcal{R}^{d\times d}$ is the parameter of LLM backbone.
@@
    where $\mathbf{\pi}_{\theta}=\{\Lambda_{t},\tilde{\Delta}_{t}\}$ and $\pi_{\text{ref}}=\Lambda_{t}$ represent the policy and reference models for round $t$ respectively, $\sigma$ is the sigmoid function, and $\beta$ is a hyper-parameter controlling the deviation from $\pi_{\text{ref}}$. Variables in the first term follow the definitions in the standard DPO loss (Rafailov et al., 2024). In the second term corresponding to cross-entropy learning towards the ground-truth video captions, $\lambda$ is the weight of the second term in the overall loss, $\mathcal{D}_{\text{gt}}$ denotes the SFT training dataset, and $(\mathbf{x},\mathbf{y}_{\text{gt}})$ corresponds to a video and its paired ground-truth text description. $\mathbf{y}_{\text{win}}$ and $\mathbf{y}_{\text{lose}}$ are preferred and dispreferred video captions generated by $\Lambda_{t}$ for $\mathbf{x}$, which are judged using a text LLM based on the atomic-event-based metric proposed in Section 3.3.1. Each mini-batch of training samples is randomly selected from $\mathcal{D}_{\text{gt}}$. It is worth noting that this gDPO loss is different from Iterative RPO (Pang et al., 2024), we use the guidance of ground truth instead of the chosen sample, which aims to stabilize training instead of avoiding probability decrease.
3.4 Transferring Caption Optimisation Gains to Video QA
Following the application of MrDPO, a significant enhancement in the model’s video captioning performance can be achieved. However, we found that the model’s general video understanding capabilities are almost entirely determined by audio-visual SFT. The subsequent RL training, which focuses exclusively on captioning, did not yield further improvements in the model’s general video understanding capabilities.
In the audio-visual SFT stage, the quality and nature of the training data play a fundamental role. Typically, this data comprises two main categories: video caption data and video QA data. We posit that video caption data is more essential of the two, especially for larger and stronger models. High-quality video captions enable the model to develop a holistic understanding of a video, encompassing its content, intricate details, underlying themes, and emotional tone. In contrast, video QA data serves to refine the model’s ability to accurately apply the knowledge it has acquired from the caption data to specific tasks.
A significant limitation of existing video caption data is that they are largely generated by existing models. While these captions can provide a general description of the video’s content, they often lack the requisite detail, which can result in models trained on this data failing to accurately perceive subtle but important aspects of the video. Consequently, we utilise the MrDPO-trained model to re-annotate existing video data, thereby creating a new, higher-quality set of video captions. We then apply this new caption data and the original QA data to directly perform audio-visual SFT after audio alignment. Through this strategy, we aim to transfer the model’s powerful video captioning performance to a more generalised and robust video understanding capability.
@@ 4.1 Model Specifications
video-SALMONN 2 series consists of video-SALMONN 2 (7B), video-SALMONN $\text{2}_{\text{F-16}}$ (7B), and video-SALMONN 2+. video-SALMONN 2 (7B) is built on an internally trained high-performance visual LLM, which is further fine-tuned on LLaVA-OneVision-7B (Li et al., 2024). The model processes video frames at a per-second frame rate of 1, and can handle up to 110 frames with $384\times 384$ pixels per frame. video-SALMONN $\text{2}_{\text{F-16}}$ (7B) is trained based on F-16 (Li et al., 2025) and processes 16 FPS video up to 1760 frames with $384\times 384$ pixels. video-SALMONN 2+ series contains 3B, 7B, and 72B versions, all derived from Qwen 2.5-VL series. The model processes video frames at 10 FPS, and can handle up to 768 frames with $61250$ pixels per frame.
For the audio branch, we use the Whisper-Large-v3 encoder (Radford et al., 2023) as the audio encoder, and a window-level Q-Former (Tang et al., 2024a) with a window length of 0.5 seconds as the audio aligner, which produces 60 audio tokens for a 30-second input. The rank $r$ and scaling factor $\alpha$ of LoRA are set to 128 and 2.0, respectively. During training, the visual encoder, the audio encoder, and the LLM remain frozen.
@@ 4.2 Data and Training Specifications
All models are trained with modality alignment. video-SALMONN 2 (7B) and video-SALMONN 2+ (7B) are trained with audio-visual SFT and MrDPO. video-SALMONN $\text{2}_{\text{F-16}}$ (7B) is trained with data annotated with video-SALMONN 2 (7B). video-SALMONN 2+ (3B) and (72B) are trained with data annotated with video-SALMONN 2+ (7B).
@@
In the audio modality alignment stage, LibriSpeech-960h (Panayotov et al., 2015) and AudioCaps (Kim et al., 2019) are used to train the audio aligner. LibriSpeech-960h is utilized for speech recognition training, while AudioCaps is employed for audio captioning training. In the audio-visual SFT stage, experiments are performed using FineVideo (Farré et al., 2024), CinePile (Rawal et al., 2024), and about 13k videos with rich audio information from LLaVA-Video-178k (Zhang et al., 2024b). Both video captioning and video QA are trained during audio-visual SFT.
@@
In the MrDPO stage, before each training round, the model generates a pair of captions for each video in the SFT dataset. To determine whether a caption pair is suitable for DPO, we evaluate the missing information rate and hallucination rate using GPT-3.5. It is also feasible to use a smaller model for evaluation, as shown in Appendix C. If deemed suitable, one caption is selected as the preferred sample, while the other is rejected. The selection criteria for each round are detailed in Appendix D. We explore the values of $\lambda$ in Eqn. (3) in Appendix E and finally set it to 0.1 throughout the entire MrDPO stage. Since only video captioning is trained during MrDPO, the video QA performance of the model will slightly decline during DPO rounds, and QA data is added in the final round. Detailed training cost can be found in Appendix B.
For further data annotation, 100k videos are randomly selected from the training datasets and are re-annotated with video-SALMONN 2 (7B) and video-SALMONN 2+ (7B).
@@
For captioning evaluation, we curated an open-sourced video captioning benchmark [^1] https://huggingface.co/datasets/videoSALMONN2/video-SALMONN_2_testset to evaluate the event missing rate (Miss) and hallucination rate (Hall). Details of the test data can be found in Appendix F. The benchmark consists of 483 carefully selected videos, each labelled with complete audio-visual captions by human annotators. Atomic events for the test dataset were initially obtained using GPT-4o and then manually refined. GPT-3.5 is used to evaluate the event missing and hallucination rates of the generated captions. Details are shown in Appendix G.
5 Experimental Results
5.1 Captioning Evaluation Results and Analysis
Table 1: Captioning evaluation and ablation on video-SALMONN 2 (7B).
(a) The comparison of captioning performance of the models. VMME denotes QA results of GPT-4o on Video-MME based on video descriptions generated by the corresponding models.
[…]
GPT-4o 17.0 14.2 31.2 46.3|2.5 64.3
Qwen2.5-VL (7B) 21.9 17.4 39.2 44.5|2.4 55.0
InternVideo 2.5 (7B) 30.8 15.0 45.8 39.6|2.2 51.8
VideoLLaMA 3 (7B) 44.9 11.6 56.5 33.4|1.9 46.3
Qwen2.5-Omni (7B) 26.7 21.7 48.1 39.7|2.2 52.7
video-SALMONN 2 (7B) 10.0 12.9 22.9 46.1|2.5 65.9
(b) The component-level ablation study conducted with video-SALMONN 2 (7B) for MrDPO.
[…]
Visual 50.7 -
+ SFT 41.8 +17.6
+ DPO 37.8 +9.6
+ gDPO 39.7 -5.0
+ LoRA Proxy 33.7 +15.1
+ MrDPO 22.9 +32.0
To demonstrate the powerful captioning capability of video-SALMONN 2 and the effectiveness of MrDPO, we evaluate video-SALMONN 2 (7B) on our caption benchmark. The performance on the "detailed" subset of the VDC caption benchmark (Chai et al., 2025) is also reported for examining visual captioning capability. Another way to evaluate caption completeness is by using the generated video captions rather than the raw video frames as input to a powerful text-based LLM for the video QA task. We used the Video-MME benchmark (Fu et al., 2025a) for testing. Specifically, we first use captioning models to generate descriptions for each video in Video-MME, and then prompt GPT-4o to answer the relevant questions based solely on these captions. In this setting, higher QA accuracy suggests that the generated captions are more complete and informative. All results are reported in Table 1(a). video-SALMONN 2 (7B) outperforms other models in both information missing and hallucination rates on our caption benchmark, as well as VDC detailed and Video-MME captioning. The high correlation for all 3 metrics shows the effectiveness of our caption benchmark.
Among existing open-source multimodal LLMs, few can provide detailed and accurate video descriptions. The caption quality of Qwen2.5-VL is the best among existing open-source models. Notably, some open-source models, such as VideoLLaMA 3, tend to generate shorter captions, leading to relatively high information missing rates but low hallucination rates. Qwen2.5-Omni can perceive both audio and visual information simultaneously, but still mainly focuses on visual information, often omitting human speech in audio and other sound events in captions. The visual-only version of GPT-4o lacks audio comprehension, leading to a higher rate of missed events. We also perform Elo ranking with human evaluators inspired by Chai et al. (2025), which validates the alignment of our proposed metrics with human perception and highlights the strong video captioning capabilities of video-SALMONN 2. Details are provided in Appendix H.
5.2 Ablation on MrDPO
To address the inherent instability of multi-round DPO, we employ two key strategies: using the gDPO loss and training with LoRA proxies. A complement component-level ablation study is shown in Table 1(b).
Specifically, Fig. 3 further demonstrates the details on two key strategies. As illustrated in Fig. 3(a), while classical DPO shows a slight advantage in the first round, the gDPO loss, which incorporates an additional cross-entropy term conditioned on ground-truth captions, consistently outperforms it from the second round onward, confirming its effectiveness in stabilizing optimization. Moreover, Fig. 3(b) shows this "Proxy" approach yields better performance than directly fine-tuning the existing LoRA ("Direct"), especially in the early stages. This improvement likely stems from the fact that merging LoRA progressively strengthens the model’s foundation by integrating prior adaptations, while re-initializing a new LoRA allows the model to flexibly explore fresh low-rank subspaces aligned with current preference signals, thus avoiding stagnation in outdated parameter spaces.
5.3 Ablation on SFT Data
Table 2: Results of models using different SFT data on Video-MME. S, M, and L denote short, medium, and long.
Model Video-MME
[…]
F-16 (7B) 65.0 78.9 63.2 52.8
+SFT data 69.2 79.8 67.1 60.7
+generated data 70.2 80.1 68.8 61.7
Qwen2.5-VL (72B) 73.3 80.8 73.5 65.4
+SFT data 78.8 82.3 81.2 72.8
+generated data 79.7 85.0 81.7 72.3
We perform audio-visual SFT on the audio-aligned model using both the original SFT data and the SFT data generated by the model after MrDPO. Specifically, we experimented with training video-SALMONN $\text{2}_{\text{F-16}}$ (7B) (Li et al., 2025) using data generated by video-SALMONN 2 (7B); and training video-SALMONN 2+ (72B) using data generated by video-SALMONN 2+ (7B). We then compared these models with those trained using the corresponding original SFT data. The evaluation results on Video-MME are shown in Table 2. Models trained with generated data consistently outperform models trained with original SFT data, which reveals that models enhanced on captioning by MrDPO can be used to generate new data and fine-tune stronger models.
5.4 Overall Results on Video QA Benchmarks
Table 3: Evaluation results of video-SALMONN 2 series models on video understanding benchmarks. Models denoted with can perceive audio-visual information, while other models can only understand visual-only frames. The best results for each scale are in bold.
Model Audio-visual QA Visual-only QA
[…]
VideoLLaMA3 (2B) 59.6 - - - - 65.4 41.6
Qwen2.5-Omni (3B) 62.0 - - - 40.5 - -
Qwen2.5-VL (3B) 61.5 - - - 37.4 68.2 43.3
video-SALMONN 2+ (3B) 68.3 48.3 66.2 42.2 67.7 70.5 48.6
video-SALMONN (13B) 43.3 - 38.3 - - - -
LLaVA-Video (7B) 63.3 40.2 56.5 - - 70.8 -
VideoLLaMA2 (7B) 54.9 25.4 44.9 - 35.2 32.7 -
VideoLLaMA3 (7B) 66.2 - - - - 73.0 45.3
InternVideo2.5 (8B) 65.1 - - - - 72.8 46.4
Qwen2.5-Omni (7B) 64.3 45.4 - 16.4 47.5 - -
Qwen2.5-VL (7B) 65.1 - - 27.8 40.7 70.2 45.3
video-SALMONN 2 (7B) 67.4 48.6 65.6 40.7 66.3 68.3 -
video-SALMONN $\text{2}_{\text{F-16}}$ (7B) 70.2 49.6 66.3 41.6 66.8 69.6 -
video-SALMONN 2+ (7B) 73.4 50.9 69.5 46.9 71.8 73.6 49.7
GPT-4o 71.9 42.6 56.6 42.0 56.5 64.6 30.8
Gemini-1.5 Pro 75.0 48.0 78.3 41.2 - - 33.1
Qwen3-Omni-Flash 71.4 54.1 - 57.3 76.2 75.5 51.1
LLaVA-Video (72B) 70.5 - - - - 74.4 -
VideoLLaMA2 (72B) 61.4 - - - - 61.2 -
Qwen2.5-VL (72B) 73.3 - - 50.2 61.8 74.6 47.3
video-SALMONN 2+ (72B) 79.7 56.5 72.2 57.8 79.4 80.4 55.5
The overall evaluation results of video-SALMONN 2 series on general video understanding benchmarks are presented in Table 3. To comprehensively evaluate the performance of video-SALMONN 2 series, we select leading models among $\sim$3B, $\sim$7B, and larger scales as our baselines. Both open-source video models, including VideoLLaMA3 (Zhang et al., 2025), LLaVA-Video (Zhang et al., 2024b), InternVideo 2.5 (Wang et al., 2025b), and Qwen2.5-VL (Bai et al., 2025), and the leading closed-source model GPT-4o are evaluated. To highlight the advantages of our model among audio-visual models, video-SALMONN (Sun et al., 2024), VideoLLaMA 2 (Cheng et al., 2024), Qwen 2.5-Omni (Xu et al., 2025a), Qwen3-Omni-Flash (Xu et al., 2025b), and Gemini-1.5 Pro (Team et al., 2024) are also compared. Audio-visual QA benchmarks include Video-MME (Fu et al., 2025a), WorldSense (Hong et al., 2025), AVUT (Yang et al., 2025b), Video-Holmes (Cheng et al., 2025), and DailyOmni (Zhou et al., 2025b). Visual-only benchmarks include MLVU (Zhou et al., 2025a) and LVBench (Wang et al., 2025a).
For audio-visual benchmarks, video-SALMONN 2+ shows dominating performance and outperforms all models with similar scales for both visual-only models and audio-visual models, even competitive to larger models. video-SALMONN 2+ (3B) outperforms all existing 7B models, and video-SALMONN 2+ (7B) is competitive to existing 72B models. The largest video-SALMONN 2+ (72B) even consistently performs better than GPT-4o and Qwen3-Omni-Flash, while surpassing Gemini-1.5 Pro on most of the audio-visual QA benchmarks. For visual-only video understanding, video-SALMONN 2+ still retains strong capabilities, achieving slight improvements compared to the base model Qwen2.5-VL, and remains highly competitive against same-scale visual-only state-of-the-art models.
6 Conclusions
We introduced video-SALMONN-2, a family of high-performance audio–visual LLMs for video captioning and video QA. Our core method, MrDPO, addresses DPO reference staleness to enable continual self-improvement, producing captions that are richer and more faithful than GPT-4o and Gemini-1.5 Pro. Using this strengthened captioner to re-annotate video data yields higher-quality supervision for SFT, and we show that gains from captioning transfer to general video understanding. Across standard video-QA benchmarks, our 3B/7B models lead their size class, and the 72B model surpasses all open-source competitors. Code, models, and data will be released to support further research.
7 Reproducibility Statement
We provide detailed descriptions of the model architecture, training steps, training settings, and datasets in Section 3 and Section 4. All non-public datasets, code, and final models will be released upon acceptance, which provides enough reproducibility for the work.
Appendix A The Use of Large Language Models
We utilized Gemini-2.5-Pro and GPT-5 to assist us in checking for grammar errors and polishing the fluency of our sentences.
Appendix B Training Cost for video-SALMONN 2 (7B)
@@
Regarding training settings, the resource consumption of video-SALMONN 2 (7B) for each training step is shown in Table 4. video-SALMONN 2+ (7B) is trained with similar settings. In our experiments, we run a total of six gDPO rounds in MrDPO.
@@
Appendix C Training with Models with Smaller Scale as Evaluator
We also try using Qwen3-4B Yang et al. (2025a) as the LLM for evaluating the quality of video captions during training, and still use GPT-3.5 for testing. All prompts, generation settings, and training settings remain the same with the GPT-3.5 version. The results of using these two models as evaluators are similar, as shown in Table 5.